# Artworks Reimagined: Exploring Human-AI Co-Creation through Body Prompting

### v1

@@ 4.4. RQ3: Participant Behavior
  • Camera booth choice: 46 participants (58%) used the installation’s public posing stage, compared to 33 participants (42%) who posed in the private booth.
@@
4.4.1. Selecting the source artwork
@@
Twelve participants mentioned they selected the artwork because it was familiar. A few participants ($n=6$) were attracted by specific objects in the source artwork. Others were attracted to the style of the artwork ($n=4$) and its mood ($n=4$). C1 mentioned memories of her childhood attracted her to the artwork, when she “lived by the lake and did laundry”. C10 mentioned being attracted to the painting because it was modern and abstract. B61, on the other hand, preferred realistic source artworks over more abstract ones.
A large proportion of the participants ($n=28$) mentioned they selected the source artwork to contrast or complement their chosen approach to body prompting. For instance, the number of subjects in the artwork was often ($n=13$) chosen to match the participants’ group size. For a few participants, the number of subjects in the artworks was perceived as a limiting factor. A24, for instance, selected an artwork with three subjects to match their group size, because : “it was the only one with three characters”.
@@
Participants often selected their source artwork prompt to match their posing strategy, aiming to either contrast (reimagine) or complement (imitate) the source artwork, as described in the following section.
4.4.2. Deciding on a body prompting strategy
When asked “Why did you pose the way you did?,” three distinct body prompting strategies emerged.
About 30% of participants body prompted to re-create the source artwork in an attempt to “mimic” (G79), “imitate” (G64), “resemble” (H11), “mirror” (F74), and “fit the artwork” (C3). Other participants simply did not want to “invent anything different” (A14). This may reflect a misunderstanding of the technology’s purpose, or simply a natural human behavior to mimic what we see displayed. However, in some cases, imitating the pose was not an easy task. E12, for instance, said she “was trying to replicate the [artwork], but the character had a leg completely up, and I did my best to replicate the pose”. Other participants also responded that their body prompt was inspired by the source artwork, such as C36 who “felt like [the pose] came from the artwork”. A13 were more specific in their imitation of the artwork: “we saw the artwork had a fishing rod. We thought lets do the same fishing rod. We were just imitating whatever was there in the artwork”. A34 responded being motivated by seeking harmony with the source artwork, “because it would adapt well to it, as it is a similar pose”. This was also reflected in E15, who said: “Well, it had to be just like that grandma pose”. B54 also mentioned being “most inspired by the artwork itself” and the pose “felt natural”.
@@ 4.4.2. Deciding on a body prompting strategy
The second posing strategy among another third of the participants was to reinterpret the artwork. Participants in this group often reported wanting to “contrast” the source artwork to create a novel image that would significantly diverge. A5, for instance, mentioned seeking “a contrast between the posing and the artwork” to force a change in narrative and create a novel artwork. A33 reported wanting “to recreate the painting in a new way”. D75 mentioned wanting “to create a different atmosphere” which hints at the underlying motivation to cause a change in narrative between the original artwork and the generated image.
@@
Finally, a third of the participants said they assumed the body prompt for no specific reason. These participants had no deep thoughts about their body prompt and simply went with their intuition in the moment, with the body prompt being “the first thing that came to mind” (A23, A24, C3) or body prompted “impulsively, on the spur of the moment” (B77), and F70 mentioned “just throwing [himself] into the moment”.
@@
4.4.3. Deciding on a specific body prompt and teamwork
@@
When asked what they wanted to express with their body prompt, participant answers were extremely diverse. Expressing affection was mentioned by a few participants who wanted to hug others (H39) or portray “friendship” (A40) and “family” (A41). Participants had fun at the event which was also reflected in their body prompts. B59, for instance, wanted to show that they were ““having fun”, and B73 mentioned she “wanted to make something silly”. For some, there was a significant acting involved in their body prompt as they assumed a pose (presumably) different from their own personality. A45, for instance, was inspired by the source artwork and wanted to express “elegance and athleticism” with a pose describing “big movements”. A35 aimed to pose as an “authoritarian evil power” trying to “scare people”. This speaks to the entertaining character and playfulness of the installation.
4.5. RQ4: Body Prompting in Public
Our installation was designed to provide a choice of body prompting in public or private environments. Almost half of the participants ($n=35$) had no strong opinion about this choice and did not reveal many concerns about privacy. A significant number of participants ($n=29$) selected to body prompt in public or private for the pragmatic reason of having a shorter wait in line. H11, for instance, selected the public booth “because it was free and the other one was occupied” while group A35 “wanted the others to see what [they were] doing” (A35). Seeing the artwork on the big public screen, as opposed to the smaller laptop screen, also motivated some participants to select the public booth.
@@ 4.5. RQ4: Body Prompting in Public
The Big Five personality factors were not found to correlate with the choice of posing area (public vs private), participation mode (group vs alone), source artwork (international vs local), dynamism of the source artwork, and expressiveness of the body prompt (Spearman’s rank correlation, $p>0.05$), with two significant, but weak exceptions. The first correlation was found to be in the personality trait of agreeableness: participants who were more agreeable tended to select a more dynamic source artwork; $\rho=0.237$, $p<0.05$, Cohen $d=0.50$ (CI [0.04, 1.01]). Second, there was a reverse correlation between extraversion and the expressiveness (dynamism) of the body prompt; $\rho=-0.236$, $p<0.05$, Cohen $d=-0.54$ (CI [-1.01, -0.10]). This indicates that for participants with high level of extraversion, the expressiveness of their body prompt tended to be low. The finding may appear counterintuitive since extraversion is typically associated with sociability, enthusiasm, and positive emotions, traits that may be associated with more dynamic and expressive poses. The inverse correlation could suggest that extraverted individuals, perhaps due to their comfort with social interactions, may not feel the need to convey their personality as strongly through a single body prompt, or they might approach body prompting with a different mindset compared to more introverted individuals. However, the correlation was very weak, as the expressiveness of the body prompt is influenced by multiple factors beyond a participant’s level of extraversion, such as the context, the individual’s mood, or the cultural background.
4.6. Other Observations
4.6.1. Key moments during the deployment
@@
4.6.2. Age appropriateness of the installation
The Researchers’ Night is an event visited by many families with children. Although children were not part of our interview study, they were allowed to use the installation in the presence of adult caretakers. This allowed an important observation: some children became very upset after using the installation. Two young children even started crying after seeing their generated image because they expected to see their likeness in the generated image. This happened even when the selected source artwork was abstract, as in the case of Jean-Michel Basquiat’s artworks. Clearly, the children had a different perception of the generated images and different expectations. For adults, the reimagined abstract artwork was interesting, but young children were often disappointed because they could not recognize themselves in the generated image.
@@ 5. Discussion
We evaluated body prompting as a novel interaction mechanism for generative AI art installations in public spaces. Overall, body prompting is a viable way of interacting with generative AI in a public setting, providing a highly engaging and fun experience for both participants and the audience. Our study identified three body prompting strategies employed by participants, focused on imitating the source artwork, re-imagining, or body prompting by “just being themselves”. Personality traits did not play an important role in the participants’ decisions and body prompts and the installation provided a novel experience which was appreciated by both extraverts and introverts.
@@ 5.1. Human-AI Co-Creation via Body Prompting
Generative AI has ushered in an era of human-AI co-creation (Kantosalo et al., 2020; Rezwana and Maher, 2023; Capel and Brereton, 2023). The human agency in this co-creative process varies on a spectrum (Parasuraman et al., 2000), from full human autonomy to full machine automation. In our study, participants experienced human-AI co-creation along this spectrum. Depending on the chosen source artwork and body prompt, the generative model had varying levels of influence on the generated image. In combination with the participants’ body prompting strategy, their experiences included either re-creation or reinterpretation, as discussed in Section 2.3. The participants pursued their strategy with varying levels of tenacity, but most interactions were casual, as participants picked a level of interaction suitable for their desired engagement (Pohl, 2015; Pohl and Murray-Smith, 2013).
While it can be argued whether the body-prompted images are art (McCormack et al., 2023; Coeckelbergh, 2023), the poietic act of performing a body prompt changed the dynamic of the creative process of image generation, placing the human in the center of the co-creative process in a public setting. In Kun et al.’s GenFrame, participants turned knobs to control image generation parameters, but otherwise relegated the creative process fully to the machine (Kun et al., 2024a). In our co-creative study, the AI became “more than just a machine”, with AI and humans being co-creators and co-performers (Coeckelbergh, 2023). This highlights the importance of interaction design in co-creative systems (Rezwana and Maher, 2023).
@@ 5.2. Recommendations for the Design of Body Prompting Installations
In this section, we reflect on the few shortcomings in the design of our study and installation, and draw on this experience to provide recommendations for practitioners interested in the design of installations and researchers conducting user studies involving interactive image generation in a GLAM context.
5.2.1. Multi-display setup and user experience design
@@
Design recommendation: Future installations should provide immediate feedback to participants without forced integration of a user study. Bottlenecks in the user flow should be avoided. We recommend to show generated images on the same screen as soon as they are available. If there is a choice between private and public body prompting, this needs to be explained at the event, for example with signage.
5.2.2. Manage expectations
As mentioned in Section 4.6.2, some participants experienced frustration. This was particular pronounced with young children who were often disappointed and did not share their parents’ appreciation for the AI-generated images. This raises a point related to managing expectations: an “AI Camera Booth” may raise expectations that the resulting images will resemble the body prompting person. Our study’s purpose was slightly different from a “camera booth”, and we carefully designed communication materials to avoid setting this expectation.
Design recommendation: Communication materials should be carefully designed to set the right expectations. The purpose of the installation should be made clear to its visitors. If what is being provided is advertised as a “camera booth”, the technology should be able to reproduce the likeness of its users. If the focus is on reinterpreting existing artworks, the term camera booth should be avoided. Facial expressions and hand gestures should be included in the body prompt.
5.2.3. Enable expressiveness
A body pose can only carry a limited amount of information. In our case, some participants reported that the generated image was more serious and sinister than expected. This could be alleviated by including the facial expression in the body prompt and providing more fine-grained control over the image generation.
Design recommendation: Include controls for users to set the mood of the generated image in addition to the body prompt, for instance with Kun et al.’s rotating knobs (Kun et al., 2024a) or other means, such as a mood selectors or sliders in the application.
5.2.4. Enhancing user experience with AI co-creativity
@@
Design recommendation: Hallucinations and weird artifacts are part of the generative AI user experience and add an intriguing component. We recommend not seeking to eliminate hallucinations, but embracing them in a playful way.
5.2.5. Interaction with the installation
To body prompt, users first needed to select a source artwork. While this interaction could take place with gestures, we decided on a touch-screen display. However, users may be unaware of interactive content (Henderson et al., 2017). We observed this interaction blindness in one case at the event where participants in the private booth faced the “first-click problem” – they did not realize the display was a touchscreen and used a mouse and keyboard that happened to be available. With the public installation, this was not an issue since participants could observe other people’s use of the installation.
Design recommendation: The first-click problem should be addressed with instructions that are carefully designed to fit the study purpose and not bias participants.. While we addressed this problem with a makeshift banner during the event, the private booth requires a slightly different design and instructions than the public posing area.
5.2.6. Takeaways and souvenirs
@@
5.2.7. Nudity and inappropriate imagery
@@ 5.2.8. Materials
5.2.9. Personality traits
Personality traits, such as extraversion and openness, were not a significant factor in our study. This is, in part, due to the specific context of the event that attracts a self-selected sample of participants with tendencies to being extraverted. The Researchers’ Night is visited by people curious to make new experiences, and our installation successfully provided these experiences. In this regard, the value of having a separate private posing area is questionable. A related issue is the nature of the private booth: it is hidden from view and easy to overlook. However, even if the booth had been better advertised, the general tendency among participants was that public booth was the preferred option, which is reflected in the visitor numbers and participant responses.
Design recommendation: We recommend not creating a private booth, and instead focus all efforts on designing the public posing area for effective engagement with the audience.

### v2

@@ 5.4.3. RQ3: Participant behavior
  • Staging area choice: 46 participants (58%) used the installation’s public posing stage, compared to 33 participants (42%) who posed in private.
@@
Selecting the source artwork
@@
Twelve participants mentioned they selected the source artwork because it was familiar to them. A few participants ($n=6$) were attracted by specific objects in the source artwork. Others were attracted to the style of the artwork ($n=4$) and its mood ($n=4$). C1 mentioned memories of her childhood attracted her to the artwork, when she “lived by the lake and did laundry”. C10 mentioned being attracted to the painting because it was modern and abstract. B61, on the other hand, preferred realistic source artworks over more abstract ones.
A large proportion of the participants ($n=28$) mentioned they selected the source artwork to either contrast or complement their chosen approach to body prompting. For instance, the number of subjects in the artwork was often ($n=13$) chosen to match the participants’ group size. For a few participants, the number of subjects in the artworks was perceived as a limiting factor. A24, for instance, selected an artwork with three subjects to match their group size, because : “it was the only one with three characters”.
@@
As mentioned above participants often selected their source artwork prompt to match their posing strategy, aiming to either contrast (reimagine) or complement (imitate) the source artwork. In the following section, we describe the different types of posing strategies.
Deciding on a body prompting strategy
When asked “Why did you pose the way you did?,” three distinct body prompting strategies emerged. About 30% of participants body prompted to re-create the source artwork in an attempt to “mimic” (G79), “imitate” (G64), “resemble” (H11), “mirror” (F74), and “fit the artwork” (C3). Other participants simply did not want to “invent anything different” (A14). This may reflect a misunderstanding of the technology’s purpose, or simply a natural human behavior to mimic what we see displayed. However, in some cases, imitating the pose was not an easy task. E12, for instance, said she “was trying to replicate the [artwork], but the character had a leg completely up, and I did my best to replicate the pose”. Other participants also responded that their body prompt was inspired by the source artwork, such as C36 who “felt like [the pose] came from the artwork”. A13 were more specific in their imitation of the artwork: “we saw the artwork had a fishing rod. We thought lets do the same fishing rod. We were just imitating whatever was there in the artwork”. A34 responded being motivated by seeking harmony with the source artwork, “because it would adapt well to it, as it is a similar pose”. This was also reflected in E15, who said: “Well, it had to be just like that grandma pose”. B54 also mentioned being “most inspired by the artwork itself” and the pose “felt natural”.
@@
The second posing strategy among another third of the participants was to reinterpret the artwork. Participants in this group often reported wanting to “contrast” the source artwork to create a novel image that would significantly diverge from the source artwork. A5, for instance, mentioned seeking “a contrast between the posing and the artwork” to force a change in narrative and create a novel artwork. A33 reported wanting “to recreate the painting in a new way”. D75 mentioned wanting “to create a different atmosphere” which hints at the underlying motivation to cause a change in narrative between the original artwork and the generated image.
@@
Finally, a third of the participants said they assumed the body prompt for no specific reason. These participants had no deep thoughts about their body prompt and simply went with their intuition in the moment, with the body prompt being “the first thing that came to mind” (A23, A24, C3) or body prompted “impulsively, on the spur of the moment” (B77). F70 mentioned “just throwing [himself] into the moment”.
@@
Deciding on a specific body prompt and teamwork
@@
When asked what they wanted to express with their body prompt, participant answers were extremely diverse. Expressing affection was mentioned by a few participants who wanted to hug others (H39) or portray “friendship” (A40) and “family” (A41). Participants had fun at the event which was also reflected in their body prompts. B59, for instance, wanted to show that they were “having fun”, and B73 mentioned she “wanted to make something silly”. For some, there was a acting involved in their body prompt as they assumed a pose (presumably) different from their own personality. A45, for instance, was inspired by the source artwork and wanted to express “elegance and athleticism” with a pose describing “big movements”. A35 aimed to pose as an “authoritarian evil power” trying to “scare people”. This speaks to the entertaining character and playfulness of the installation.
5.4.4. RQ4: Body prompting in public
Our installation was designed to provide a choice of body prompting in a public or private setting. Almost half of the participants ($n=35$) had no strong opinion about this choice and did not reveal concerns about privacy. A significant number of participants ($n=29$) selected to body prompt in public or private for the pragmatic reason of having a shorter wait in line. H11, for instance, selected the public booth “because it was free and the other one was occupied” while group A35 “wanted the others to see what [they were] doing” (A35). Seeing the artwork on the big public screen, as opposed to the smaller laptop screen, also motivated some participants to select the public booth.
@@ 5.4.4. RQ4: Body prompting in public
The Big Five personality factors were found to have no significant correlation with the choice of posing area (public vs private), participation mode (group vs alone), source artwork (international vs local), dynamism of the source artwork, and expressiveness of the body prompt (Spearman’s rank correlation, $p>0.05$), with two significant, but weak exceptions. The first correlation was found to be in the personality trait of agreeableness: participants who were more agreeable tended to select a more dynamic source artwork; $\rho=0.237$, $p<0.05$, Cohen $d=0.50$, 95% CI = [0.04, 1.01]. Second, there was a reverse correlation between extraversion and the expressiveness (dynamism) of the body prompt; $\rho=-0.236$, $p<0.05$, Cohen $d=-0.54$, 95% CI = [-1.01, -0.10]. This indicates that for participants with high level of extraversion, the expressiveness of their body prompt tended to be low. The finding may appear counterintuitive since extraversion is typically associated with sociability, enthusiasm, and positive emotions, traits that may be associated with more dynamic and expressive poses. The inverse correlation could suggest that extraverted individuals, perhaps due to their comfort with social interactions, may not feel the need to convey their personality as strongly through a single body prompt, or they may approach body prompting with a different mindset compared to more introverted individuals. However, the correlation was very weak, as the expressiveness of the body prompt is influenced by multiple factors beyond a participant’s level of extraversion, such as the context, the individual’s mood in-the-moment, or the cultural background.
5.4.5. Other observations
Key moments during the deployment
@@ 5.4.5. Other observations
Age appropriateness of the installation
The Researchers’ Night is an event visited by many families with children. Although children were not part of our interview study, they were allowed to use the installation in the presence of adult caretakers. This allowed an important observation: some children became very upset after using the installation. Two young children started crying after seeing their generated image because they expected to see their likeness in the generated image. Clearly, the generative system had an emotional impact on some children, conjuring up negative emotions (see also Alessandro et al. (2024)). This happened even when the selected source artwork was abstract, as in the case of Jean-Michel Basquiat’s artworks. Clearly, the children had different expectations than adults and a different appreciation for the generated image. For adults, the reimagined abstract artwork was interesting, but young children were often disappointed because they could not recognize themselves in the generated image.
@@ 6. Discussion
We evaluated body prompting as a novel interaction mechanism for generative AI art installations in a public event setting. Overall, body prompting is a viable way of interacting with image generation AI, providing a highly engaging and fun experience for both the active users and on-looking audience. Our study identified three body prompting strategies employed by participants, focused on imitating the source artwork, re-imagining, or body prompting by “just being themselves”. Personality traits did not play an important role in the participants’ decisions and body prompts and the installation provided a novel experience which was appreciated by both extraverts and introverts.
@@ 6.1. Human-AI Co-Creation via Body Prompting
Generative AI has ushered in an era of human-AI co-creation (Kantosalo et al., 2020; Rezwana and Maher, 2023; Capel and Brereton, 2023). The human agency in this co-creative process varies on a spectrum (Parasuraman et al., 2000), from full human autonomy to full machine automation. In our study, participants experienced human-AI co-creation along this spectrum. Depending on the seleced source artwork and body prompt strategy, the generative AI had varying levels of influence on the generated image. In combination with the participants’ body prompting strategy, the participants’ experiences included either re-creation or reinterpretation, as discussed in Section 2.3. The participants pursued their body prompting strategy with varying levels of tenacity, but most interactions were casual, as participants picked a level of interaction suitable for their desired engagement (Pohl, 2015; Pohl and Murray-Smith, 2013).
While it can be argued whether the body-prompted images are art (McCormack et al., 2023; Coeckelbergh, 2023), the poietic act of performing a body prompt changed the dynamic of the creative process of image generation, placing the human in the center of the co-creative process in a public setting. In Kun et al.’s GenFrame, participants turned knobs to control image generation parameters, but otherwise relegated the creative process fully to the machine (Kun et al., 2024a). In our co-creative study, the AI became “more than just a machine”, with AI and humans being co-creators and co-performers (Coeckelbergh, 2023). This highlights the importance of interaction design in the design of co-creative systems (Rezwana and Maher, 2023).
6.2. Ethical Considerations
In designing and conducting our study, we considered the ethical implications and potential harms that may emerge due to taking pictures of participants in public. This section elucidates the steps taken to mitigate potential harm.
6.2.1. Consent and data privacy
Public photography involves capturing people’s visual identities, potentially infringing on their privacy rights. While it is legally permissible in many places to photograph individuals in public places without consent, this can still raise concerns, especially when individuals are the central focus of the photograph. Although public settings do not guarantee privacy, individuals expect not to be intentionally photographed for studies without their consent. To respect the participants’ privacy, we implemented anonymization protocols by storing only the detected pose, not the raw photograph. Participants were informed on the consent screen about data collection, the study’s purpose, the non-storage of original photos, and their right to withdraw at any time without penalty. Further, photographic data, especially when linked to identities, are sensitive in nature. We ensured the use of robust security measures to protect the data from unauthorized access, and we committed to using the collected data solely for the purposes of this study.
6.2.2. Body shaming and self-esteem
Photographic representations can sometimes lead to harmful body comparisons or negative self-perception, especially if individuals are self-conscious about their appearance. We strived to conduct our study in a manner that respects the dignity, privacy, and emotional well-being of our participants. We took particular care to avoid any forms of body shaming or implicit biases in our study design and data interpretation. To this end, we only collected the pose of participants (detected with openpose), rather than the original photo. The presentation of results was, therefore, neutral, focusing on behavioral insights rather than personal attributes. The pose information was coarse, without identifying facial expressions or details.
6.2.3. Emotional and psychological impact
We acknowledge that viewing one’s own image unexpectedly, as was the case in aour installation where the user saw a webcam video stream during the countdown, might result in unexpected emotional or psychological reactions in some users. Further, to minimize potential distress, we implemented a debriefing process where participants were given the opportunity to discuss their feelings about the photographs and the study in general. This process also served as a platform to offer support and resources, should participants feel negatively impacted.
@@ 6.3. Recommendations for the Design of Body Prompting Installations
In this section, we reflect on the few shortcomings in the design of our study and installation, and draw on this experience and the interviews with participants to provide recommendations for practitioners interested in the design of installations and for researchers conducting user studies involving interactive image generation in the context of GLAM institutions or public event settings.
Multi-display setup and user experience design
@@
Design recommendation: Future installations should provide immediate feedback to participants without forced integration of a user study. Bottlenecks in the user flow should be avoided. We recommend showing the resulting images on the same screen as soon as they are available. If there is a choice between private and public body prompting, this needs to be explained at the event, for example with signage.
Manage expectations
As mentioned in Section 5.4.5, some participants experienced frustration. This was particular pronounced with young children who were often disappointed and did not share their parents’ appreciation for the AI-generated images. This raises a point related to managing expectations: an “AI Camera Booth” may raise expectations that the resulting images will closely resemble the body prompting person. Our study’s purpose was slightly different from a “camera booth”, and we carefully designed our communication materials to avoid setting this expectation.
Design recommendation: The purpose of the installation should be made clear to its visitors. Communication materials should be carefully designed to set the right expectations. If what is being provided is advertised as a “camera booth”, the technology should be able to reproduce the likeness of its users. If the focus is on reinterpreting (or reimagining) existing artworks, the term camera booth should be avoided. Facial expressions and hand gestures should be included in the body prompt.
Enable expressiveness
A body prompt can only carry a limited amount of information. In our case, some participants reported that the generated image was more serious and sinister than expected. This could be alleviated by including the facial expression in the body prompt and providing more fine-grained control over the image generation.
Design recommendation: Include controls for users to set the mood of the generated image in addition to the body prompt, for instance with Kun et al. (2024a)’s rotating knobs or other means, such as a mood selectors or sliders in the application.
Enhancing user experience with AI co-creativity
@@
Design recommendation: Hallucinations and weird artifacts are part of the generative AI user experience and add an intriguing component. We recommend not seeking to eliminate hallucinations entirely. Instead, the the system should embrace hallucinations and surprising interactions in a playful way.
Interaction with the installation
In our installation, in order to body prompt, users first needed to select a source artwork. While this interaction could take place with gestures, we decided on a touch-screen display. However, users may be unaware of interactive content (Henderson et al., 2017). We observed this interaction blindness in one case at the event where participants in the private booth faced the “first-click problem” – they did not realize the display was a touchscreen and used a mouse and keyboard that happened to be available. With the public installation, this was not an issue since participants could observe other participants’ use of the installation.
Design recommendation: The first-click problem should be addressed with instructions that are carefully designed to fit the study purpose and not bias participants. While we addressed this problem with a makeshift banner during the event, the private booth requires a slightly different design and instructions than the public posing stage.
Takeaways and souvenirs
@@
Nudity and inappropriate imagery
@@
Personality traits
Personality traits, such as extraversion and openness, were not a significant factor in our study. This is, in part, due to the specific context of the event that attracts a self-selected sample of participants with tendencies to being extraverted. The Researchers’ Night is visited by people curious to make new experiences, and our installation successfully provided these experiences. In this regard, the value of having a separate private posing booth is questionable. A related issue is the nature of the private booth: it is hidden from view and easy to overlook. However, even if the booth had been better advertised, the general tendency among participants was that public booth was the preferred option, which is reflected in the visitor numbers and participant responses.
Design recommendation: We recommend not creating a private booth, and instead focus all efforts on designing the public posing stage for effective engagement with the audience.